OpenStudy (anonymous):

A spring in, when not taut, has length 10cm. suppose it takes 1kg m/sec^2 to compress the spring 1cm. What is the work required to stretch the spring 3cm?

OpenStudy (amistre64):

W = F*d ; We integrate this somehow, gonna have to refresh meself on it tho

OpenStudy (amistre64):

{S} f(x) dx where f(x) = hookes law; kx

OpenStudy (amistre64):

k = 1kg and x = 1 cm; so ....still reading up on it lol

OpenStudy (amistre64):

oy its been awhile ..... 1000 kg m/sec^2 = k(.01m) maybe?

OpenStudy (amistre64):

k = 100 kg/sec^2 then i think

OpenStudy (amistre64):

the int this 100x (kg m/sec^2) from 0 to .03 m

OpenStudy (amistre64):

100x^2 ------- = 50x^2 at x=.03 2 50*.0009 = .045 maybe... I aint got the textbook with me to be sure :)

OpenStudy (anonymous):

i will make sure to double check with the t/b thanks amistre64 you are a big help! :)

OpenStudy (amistre64):

youre welcome :) it sounds right, but I just cant be sure ...its been to long since I tried doing spring calculus lol